TheerealShawnteee:

What is the energy equivalent of an object with a mass of 1.83 kg? 1.65 × 1017 J 2.92 × 1017 J 3.10 × 1017 J 5.49 × 1017 J i need help in eath science pls

AZ:

e = mc^2 e is your energy m is your mass c is the speed of light which is a constant 3.0 * 10^8 m/s^2 so we know m = 1.83 kg so plug it inL e = 1.83 * (3.0*10^8)^2 e = ??

TheerealShawnteee:

um huh

TheerealShawnteee:

i dont understand im dumb im sorry

AZ:

do you have a scientific calculator?

TheerealShawnteee:

no but google does

AZ:

yes, so did you plug it in to google's calculator?

TheerealShawnteee:

plug wht tho like does the question give us enough info im so lost

AZ:

Your question is What is the energy equivalent of an object with a mass of 1.83 kg? Can you read that out loud to yourself?

AZ:

They want to calculate the ENERGY of an object that has a MASS of 1.83 kilograms

AZ:

There's this formula, you've probably heard it- Einstein came up with it E equals to m c squared E = mass * (speed of light)^2 or E = mc^2 E is energy m is mass c is speed of light the speed of light is a constant (approximately 3*10^8 m/s^2

AZ:

Since the speed of light is a constant, all you do is multiply c^2 with the mass and you'll get the energy

AZ:

Any of that make sense? @TheerealShawnteee

TheerealShawnteee:

yes

TheerealShawnteee:

@AZ it does ty

AZ:

np!

AZ:

Let me know what you get as your final answer- I'd be happy to check it for you :)